OpenStudy (abb0t):

Find the term independent of x in the expansion of:

OpenStudy (abb0t):

\[(2x+\frac{ 1 }{ x })^{10}\]

OpenStudy (anonymous):

u know binomial expansion.........?

OpenStudy (anonymous):

for term independent of x the power of x should be 0

OpenStudy (anonymous):

general term in the expansion of (2x + 1/x)^10 is C(10,r )*(2x)^(10-r)*(x)^(-r)=C(10,r )*(2)^(10-r) *x^(10-2r)

OpenStudy (abb0t):

hwat? Lol

OpenStudy (anonymous):

for the term independent of x 10-2r =0 r=5 hence (5+1)th or 6th term is independent of x

OpenStudy (anonymous):

its coefficient is given by C(10,5) * 2^5

OpenStudy (abb0t):

@nitz r u suggesting that I expand this to the 10th power?! Why?! WHY?!

OpenStudy (anonymous):

no i didnt say that...

OpenStudy (anonymous):

i just asked u that u know binomial as this question belongs to that topic..

OpenStudy (abb0t):

oh.